Shift report 20030215 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sat Feb 15 2003 - 19:18:02 EST

  • Next message: Apache: "Shift report 20030215 16:00-24:00"
    ------------------------------------------------------------
    Date: Sat Feb 15 19:18:02 2003, Shift: 08:00-16:00
    Supervisor: JHL, with: 
    Log book#14 pages 19 - 19
    
    Shift Summary:
    ------------------------------------------------------------
    Beam: ~4 Hrs Beam starting ZDC rate ~ 4500/sec.
    
    Data: In the process of taking 
          FS 12deg 1/4 200K
          MRS 90deg 6KB 500K
    
    Problems:
      -Multiple power trips in Rack4 (and Rack3,5)
       DAQ problems caused by the power trips.
      (See e-log for details)
      -T3 lost signals in some planes. The problems
       in the runs from yesterdays also. All HV's are
       fine. (See e-log)   
      -H2 still has no signal in 2 channels.  
    
    Plan:
      We have almost finished the initial low-pt scan.
      MRS has been taking high-pt data at 90 deg.
      (So far ~2M Trigger3 for 6B and ~1.5M Trigger3 for 6A)
      For FS, after we finish 12 1/4A 200K, we will start
      to collect High-pt data starting from 12 deg.
    
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Sat Feb 15 2003 - 19:21:34 EST